# H-STAR: LLM-driven Hybrid SQL-Text Adaptive Reasoning on Tables

### v2

@@ Abstract
Tabular reasoning involves interpreting natural language queries about tabular data, which presents a unique challenge of combining language understanding with structured data analysis. Existing methods employ either textual reasoning, which excels in semantic interpretation but struggles with mathematical operations, or symbolic reasoning, which handles computations well but lacks semantic understanding. This paper introduces a novel algorithm H-STAR that integrates both symbolic and semantic (textual) approaches in a two-stage process to address these limitations. H-STAR employs: (1) step-wise table extraction using ‘multi-view’ column retrieval followed by row extraction, and (2) adaptive reasoning that adapts reasoning strategies based on question types, utilizing semantic reasoning for direct lookup and complex lexical queries while augmenting textual reasoning with symbolic reasoning support for quantitative and logical tasks. Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency. [^1] The codes and data are available at: https://github.com/nikhilsab/H-STAR
@@ 1 Introduction
Tabular data is one of the most widely used formats for storing structured information in real-world applications. Table-based reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al. (2023). Thus, understanding and inferring tabular data has become a significant area of research in natural language processing. Tabular reasoning tasks (illustrated in Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al. (2022) and table-based fact verification Chen et al. (2019), have been extensively explored in the past.
Figure 1: An illustration of different tabular reasoning tasks (a) Fact-verification, (b) Short-form QA, and (c) Long-form QA. For each task, a question Q is paired with its answer A, which varies by task. Evidence shows the relevant columns and rows needed to answer the question.
@@
Advancements in Large Language Models (LLMs) have led to better performance across various tasks through carefully crafted prompts. In the domain of table reasoning, symbolic approaches such as Program of Thought Chen et al. (2023) and textual methods like Chain of Thought (CoT) Wei et al. (2022); Brown et al. (2020) have been explored. However, when applied individually (see Figure 2) these methods often struggle due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al. (2023); Liu et al. (2023); Sui et al. (2024). Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning. Conversely, SQL-based approaches are strong in quantitative problem-solving but perform poorly on noisy or unstructured inputs Liu et al. (2023). As a result, recent techniques Cheng et al. (2022); Ye et al. (2023); Wang et al. (2023); Nahid and Rafiei (2024) that rely solely on textual or symbolic reasoning suffer from similar limitations. Therefore, it is crucial to explore the question: How can symbolic and textual approaches be integrated into a hybrid method for tabular reasoning that leverages their complementary strengths while mitigating their individual limitations?
@@
In this paper, we introduce H-STAR which combines symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds. H-STAR decomposes the table reasoning task into two sub-tasks: (a) Table Extraction and (b) Adaptive Reasoning. In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using the original table and its transposed form. The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection Ye et al. (2023); Nahid and Rafiei (2024). Focusing on relevant table cells helps LLMs by providing the right context, thereby reducing hallucinations in reasoning. We employ a combination of SQL-based and text-based techniques in H-STAR’s table extraction phase. In the adaptive reasoning phase, we use the language comprehension capabilities of LLMs to guide them with few-shot examples to decide when to support semantic methods with symbolic (SQL) approaches. H-STAR employs semantic reasoning universally, using it exclusively for direct lookup, common-sense, and complex lexical queries while using an additional SQL step for quantitative, mathematical, and logical tasks. This strategy optimizes performance on diverse question types, supporting text comprehension with computation when required.
@@
We demonstrate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification. Our experiments show that using the combined approach of H-STAR, aligned with multi-view extraction and adaptive reasoning outperforms all prior state-of-the-art approaches. Our analysis demonstrates that H-STAR is robust and generalizable across LLMs, showcasing its superior reasoning capabilities across various tasks. Our main contributions are:
@@
(b) Through extensive experiments, we demonstrate that H-STAR’s hybrid approach outperforms individually implemented symbolic and semantic methods, particularly excelling on longer tables.
(c) Our experiments showcase H-STAR’s efficient table extraction and enhanced reasoning capabilities across different models and datasets, emphasizing the importance of every stage in achieving improved performance.
@@ 2 H-STAR Approach
H-STAR decomposes the table reasoning task into extraction and reasoning stages, combining LLM’s textual comprehension with symbolic reasoning. It converts the original table (T) into a query-specific table (TCR) (refer Algorithm 1). Figure 3 illustrates our H-STAR approach. Unlike, DATER Ye et al. (2023) and TabQSLify Nahid and Rafiei (2024), which use textual reasoning and text-to-SQL respectively, H-STAR integrates both reasoning types in a complementary manner for table extraction. Our reasoning step dynamically uses symbolic techniques alongside semantic methods for quantitative questions, overcoming the limits of pure textual reasoning. H-STAR operates in two main stages: 1) Table Extraction and 2) Adaptive Reasoning.
@@ 2.1 Table Extraction
Figure 3: An overview of our proposed method, H-STAR, consisting of a combination of code generation and text-based verification. Given a complex table and its question, H-STAR answers using (a) Table Extraction: extracts the question-specific table from the original by first selecting the columns followed by rows. (b) Adaptive Reasoning: when the question has any mathematical component, it generates an additional table using SQL used in the textual reasoning step.
@@
As shown in Algorithm 1, H-STAR uses a ‘multi-view’ technique, first using the original table T (line 1) followed by its transposed form TT (line 2) for column extraction. Changing the ‘view’ acts as a verifier accounting for any information missed due to the table structure, thus improving the overall performance (see Appendix A.4). In Figure 3, given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate an SQL query to extract columns ‘year’ (C1). This is followed by the text-based step coltext which processes table TT to return columns C2 (‘year’, ‘national cup’). This step captures the ‘national cup’ information relevant to answer the query, which is missed by colsql. Table T is then filtered for columns ‘year’, ‘national cup’ (C’) derived from C1 and C2 to obtain table Tc (lines 3,4).
@@
After obtaining the filtered table TC with relevant columns (Algorithm 1, line 4), H-STAR proceeds to the row extraction phase culminating in the query-specific table TCR (lines 5-8). The filtered table TC has fewer tokens than T, fitting within the token limit and enabling more efficient row extraction. As shown in Figure 3, the SQL query generated by rowsql on the column-filtered table TC extracts the relevant rows R1 (‘row 18’). While it returns the appropriate row, it is not sufficient to answer the question. To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ (R2). Combining R1 and R2 gives the relevant rows R’ (‘row 1’, ‘row 18’), resulting in the final table TCR.
@@ 2.2 Adaptive Reasoning
We propose an adaptive reasoning framework that harnesses the strengths of textual reasoning while mitigating its quantitative limitations through symbolic reasoning. First, we prompt LLMs to leverage their language understanding capabilities to evaluate the query requirements. We apply symbolic reasoning for queries necessitating quantitative analysis and use the generated output (Ev) along with the table (TCR) as the input to the final textual reasoning step. For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM needs to calculate the answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code. This code is executed on a SQL engine, enabling precise answers from the table TCR. The output from this SQL-based evaluation Ev serves as additional evidence, supporting the final reasoning step (ftext). By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling quantitative reasoning, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
@@ 3 Experiments
We evaluate our method on three datasets covering fact verification, short-form, and long-form question-answering tasks using in-context examples. (a) TabFact Chen et al. (2019): A fact verification benchmark utilizing Wikipedia tables. We evaluate the test-small set, containing 2,024 statements and 298 tables, with each statement labeled as Entailed ("True") or Refuted ("False"). (b) WikiTQ Pasupat and Liang (2015): The WikiTableQuestions (WikiTQ) dataset involves question-answering tasks over semi-structured Wikipedia tables. It includes a standard test set with 4,344 table-question pairs. (c) FeTaQA Nan et al. (2022): FeTaQA (Free-Form Table Question Answering) comprises free-form questions requiring synthesis from multiple table sections for responses. It demands advanced reasoning and profound tabular data understanding, evaluated on a test set of 2003 samples.
@@
We tailor our evaluation metrics based on the task and dataset characteristics. For fact-verification datasets like TabFact, we evaluate performance using binary classification accuracy. For short-form question-answering datasets such as WikiTQ, we assess accuracy by measuring the exact match between predicted outputs and the gold-standard answers. For more complex tasks in FeTaQA, which involve long-form question answering, we evaluate performance using ROUGE-1, ROUGE-2, and ROUGE-L metrics Lin (2004), comparing predicted outputs against the long-form answers.
@@
(a) Generic Reasoning. These methods direct the LLM to carry out the required downstream task based on the information from the table and the input question. They include End-to-End QA, which offers only the table and the question, and Few-Shot QA, which involves a few examples with the table-question-answer triplet alongside the table and the question. Chain-of-Thought prompts the LLMs to provide a supporting reasoning chain that leads to the answer.
(b) Table Manipulation. These techniques involve several steps, with the initial stage dedicated to automatically pre-processing the table for the reasoning task. BINDER utilizes in-context examples to produce SQL or Python code containing LLM API calls to generate answers. DATER directs the language model to break down tables and questions, applying reasoning to the decomposed tables based on sub-questions. Chain-of-Table employs the table as an intermediate output in its reasoning process, iterating through tasks until the final answer is obtained. TabSQLify uses SQL to trim the table and then reasons based on the reduced table.
@@ 3.1 Main Results
Table 1 compares the performance of different methods on the TabFact, and WikiTQ datasets, across GPT-3.5-Turbo and PaLM-2. This comparison involves evaluating against generic reasoning, table manipulation techniques, and H-STAR. Appendix A includes analyses on the FeTaQA dataset.
@@
Comparison Across Methods
Table 2 compares H-STAR with ReAcTable Zhang et al. (2023) and SYNTQA Zhang et al. (2024b), which use both textual and symbolic approaches for table reasoning. SYNTQA (GPT) uses the GPT model to decide between semantic and symbolic approaches, applying the chosen method for reasoning based on its selection. Additionally, it is compared with ALTER Zhang et al. (2024a), which augments both queries and table data to facilitate reasoning.
@@
Analysis. H-STAR outperforms methods that use symbolic and textual approaches, such as ReAcTable, SYNTQA(GPT), and ALTER, showing the effectiveness of its complementary reasoning in a two-stage process. Appendix A.1 presents a comparison with more methods.
Performance Across LLMs
@@ 3.2 Efficiency Analysis
Analysis. In particular, H-STAR significantly reduces the average number of processed table cells - from 159 to 18 for the WikiTQ dataset, from 88 to 13 for the TabFact dataset, and from 86 to 10 for the FeTaQA dataset. In contrast, other table manipulation methods, like TabSQLify and DATER yield higher values. Our approach showcases superior effectiveness, notably reducing cell counts across all datasets. This emphasizes how H-STAR effectively reduces irrelevant information, thus focusing LLMs on the right evidence for right reasoning.

### v3

@@ Abstract
Tabular reasoning involves interpreting natural language queries about tabular data, which presents a unique challenge of combining language understanding with structured data analysis. Existing methods employ either textual reasoning, which excels in semantic interpretation but struggles with mathematical operations, or symbolic reasoning, which handles computations well but lacks semantic understanding. This paper introduces a novel algorithm H-STAR that integrates both symbolic and semantic (textual) approaches in a two-stage process to address these limitations. H-STAR employs: (1) step-wise table extraction using ‘multi-view’ column retrieval followed by row extraction, and (2) adaptive reasoning that adapts reasoning strategies based on question types, utilizing semantic reasoning for direct lookup and complex lexical queries while augmenting textual reasoning with symbolic reasoning support for quantitative and logical tasks. Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency. [^1] The code and data are available at: https://github.com/nikhilsab/H-STAR
@@ 1 Introduction
Tabular data are among the most widely used formats for storing structured information in real-world applications. Tabular reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al. (2023). Thus, understanding and inferring tabular data has become a significant area of research in natural language processing. Tabular reasoning tasks (Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al. (2022) and table-based fact verification Chen et al. (2019), have been extensively explored in the past.
Figure 1: An illustration of different tabular reasoning tasks (a) Fact-verification, (b) Short-form QA, and (c) Long-form QA. For each task, question Q is paired with its answer A, which varies by task. Evidence shows the relevant columns and rows needed to answer the question.
@@
Advancements in Large Language Models (LLMs) have led to better performance across various tasks through carefully crafted prompts. In the domain of table reasoning, symbolic approaches such as Program of Thought Chen et al. (2023) and textual methods such as Chain of Thought (CoT) Wei et al. (2022); Brown et al. (2020) have been explored. However, when applied individually (see Figure 2) these methods often struggle due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al. (2023); Liu et al. (2023); Sui et al. (2024). Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning. In contrast, SQL-based approaches are strong in quantitative problem solving, but perform poorly on noisy or unstructured inputs Liu et al. (2023). As a result, recent techniques Cheng et al. (2022); Ye et al. (2023); Wang et al. (2023); Nahid and Rafiei (2024) that rely solely on textual or symbolic reasoning suffer from similar limitations. Therefore, it is crucial to explore the question: How can symbolic and textual approaches be integrated into a hybrid method for tabular reasoning that leverages their complementary strengths while mitigating their individual limitations?
@@
In this paper, we introduce H-STAR which combines symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds. H-STAR decomposes the table reasoning task into two sub-tasks: (a) Table Extraction and (b) Adaptive Reasoning. In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using the original table and its transposed form. The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection Ye et al. (2023); Nahid and Rafiei (2024). Focusing on relevant table cells helps LLMs by providing the right context, thus reducing hallucinations in reasoning. We employ a combination of SQL-based and text-based techniques in H-STAR’s table extraction phase. In the adaptive reasoning phase, we use the language comprehension capabilities of LLMs to guide them with few-shot examples to decide when to support semantic methods with symbolic (SQL) approaches. H-STAR employs semantic reasoning universally, using it exclusively for direct lookup, common sense, and complex lexical queries while using an additional SQL step for quantitative, mathematical, and logical tasks. This strategy optimizes performance on diverse question types, supporting text comprehension with computation when required.
@@
We demonstrate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification. Our experiments show that using the combined approach of H-STAR, aligned with multi-view extraction and adaptive reasoning outperforms all prior state-of-the-art approaches. Our analysis demonstrates that H-STAR is robust and generalizable across LLMs, showcasing its superior reasoning capabilities across various tasks. Our main contributions are as follows:
@@
(b) Through extensive experiments, we show that H-STAR’s hybrid approach surpasses state-of-the-art methods that rely solely on either symbolic or semantic techniques, especially when reasoning over larger tables.
(c) Our ablation study and analysis highlight H-STAR’s effectiveness in table extraction and reasoning across various models and datasets, underscoring the significance of each stage in enhancing overall performance.
The code and additional resources are available at: https://hstar-llm.github.io/.
@@ 2 H-STAR Approach
H-STAR decomposes the table reasoning task into extraction and reasoning stages, combining LLM’s textual comprehension with symbolic reasoning. It converts the original table (T) into a query-specific table (TCR) (refer Algorithm 1). Figure 3 illustrates our H-STAR approach. Unlike DATER Ye et al. (2023) and TabQSLify Nahid and Rafiei (2024), which use textual reasoning and text-to-SQL respectively, H-STAR integrates both reasoning types in a complementary manner for table extraction. Our reasoning step dynamically uses symbolic techniques alongside semantic methods for quantitative questions, overcoming the limits of pure textual reasoning. H-STAR operates in two main stages: 1) Table Extraction and 2) Adaptive Reasoning. Implementation details, including prompts, input table formats, and hyperparameters, are provided in Appendix C and D.
@@ 2.1 Table Extraction
Figure 3: An overview of H-STAR, consisting of a combination of code generation and text-based verification. Given a complex table and its question, H-STAR answers using (a) Table Extraction: extracts the question-specific table from the original by first selecting the columns followed by rows. (b) Adaptive Reasoning: when the question has any mathematical component, it generates an additional table using SQL used in the textual reasoning step.
@@
As shown in Algorithm 1, H-STAR uses a ‘multi-view’ technique, first using the original table T (line 1) followed by its transposed form TT (line 2) for column extraction. Changing the ‘view’ acts as a verifier accounting for any information missed due to the table structure, thus improving the overall performance (see Appendix A.3). In Figure 3, given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate an SQL query to extract columns ‘year’ (C1). This is followed by the text-based step coltext which processes table TT to return columns C2 (‘year’, ‘national cup’). This step captures the ‘national cup’ information relevant to answering the query, which is missed by colsql. Table T is then filtered for columns ‘year’, ‘national cup’ (C’) derived from C1 and C2 to obtain table Tc (lines 3,4).
@@
After obtaining the filtered table TC with relevant columns (Algorithm 1, line 4), H-STAR proceeds to the row extraction phase culminating in the query-specific table TCR (lines 5-8). The filtered table TC has fewer tokens than T, fitting within the token limit and enabling more efficient row extraction. As shown in Figure 3, the SQL query generated by rowsql on the column-filtered table TC extracts the relevant rows R1 (‘row 18’). Although it returns the appropriate row, it is not sufficient to answer the question. To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ (R2). Combining R1 and R2 gives the relevant rows R’ (‘row 1’, ‘row 18’), resulting in the final table TCR.
@@ 2.2 Adaptive Reasoning
We propose an adaptive reasoning framework that harnesses the strengths of textual reasoning while mitigating its quantitative limitations through symbolic reasoning. First, we prompt LLMs to leverage their language understanding capabilities to evaluate the query requirements. We apply symbolic reasoning for queries necessitating quantitative analysis and use the generated output (Ev) along with the table (TCR) as the input to the final textual reasoning step. For instance, in Figure 3, when faced with a question like ‘How long did it take the New York Americans to win the National Cup after 1936?’, LLM needs to calculate the answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code. This code is executed on a SQL engine, enabling precise answers from the table TCR. The output from this SQL-based evaluation Ev serves as additional evidence, supporting the final reasoning step (ftext). By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling quantitative reasoning, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
@@ 3 Experiments
We evaluate our method on three datasets covering fact verification, short-form, and long-form question-answering tasks using in-context examples. (a) TabFact Chen et al. (2019): A fact verification benchmark utilizing Wikipedia tables. We evaluated the test-small set, containing 2,024 statements and 298 tables, with each statement labeled Entailed ("True") or Refuted ("False"). (b) WikiTQ Pasupat and Liang (2015): The WikiTableQuestions (WikiTQ) dataset involves question-answering tasks over semi-structured Wikipedia tables. It includes a standard test set with 4,344 table-question pairs. (c) FeTaQA Nan et al. (2022): FeTaQA (Free-Form Table Question Answering) comprises free-form questions requiring synthesis from multiple table sections for responses. It demands advanced reasoning and profound tabular data understanding, evaluated on a test set of 2003 samples.
@@
We tailor our evaluation metrics based on the task and dataset characteristics. For fact verification data sets such as TabFact, we evaluated performance using binary classification accuracy. For short-form question-answering datasets such as WikiTQ, we assess accuracy by measuring the exact match between predicted outputs and the gold-standard answers. For more complex tasks in FeTaQA, which involve long-form question answering, we evaluate performance using ROUGE-1, ROUGE-2, and ROUGE-L metrics Lin (2004), comparing predicted outputs with long-form answers.
@@
(a) Generic Reasoning. These methods direct the LLM to carry out the required downstream task based on the information from the table and the input question. They include End-to-End QA, which offers only the table and the question, and Few-Shot QA, which involves a few examples with the table-question-answer triplet alongside the table and the question. Chain-of-Thought prompts the LLMs to provide a supporting reasoning chain that leads to the answer. (b) Table Manipulation. These techniques involve several steps, with the initial stage dedicated to automatically pre-processing the table for the reasoning task. BINDER utilizes in-context examples to produce SQL or Python code containing LLM API calls to generate answers. DATER directs the language model to break down tables and questions, applying reasoning to the decomposed tables based on sub-questions. Chain-of-Table employs the table as an intermediate output in its reasoning process, iterating through tasks until the final answer is obtained. TabSQLify uses SQL to trim the table and then reasons based on the reduced table.
@@ 3.1 Main Results
Table 1 compares the performance of different methods on the TabFact, and WikiTQ datasets, across GPT-3.5-Turbo and PaLM-2. This comparison involves evaluating against generic reasoning, table manipulation techniques, and H-STAR. Appendix A.1 offers a comprehensive analysis of the FeTaQA dataset, including a comparison with baselines, human evaluation, and qualitative analysis.
@@
Comparison Across Methods.
Table 2 compares H-STAR with ReAcTable Zhang et al. (2023) and SYNTQA Zhang et al. (2024b), which use both textual and symbolic approaches for table reasoning. SYNTQA (GPT) uses the GPT model to decide between semantic and symbolic approaches, applying the chosen method for reasoning based on its selection. Additionally, it is compared with ALTER Zhang et al. (2024a), which augments both queries and table data to facilitate reasoning and E5 Zhang et al. (2024c), leveraging code generation capabilities of LLMs for table extraction followed by reasoning.
@@
Analysis. H-STAR outperforms methods that use symbolic and textual approaches, such as ReAcTable, SYNTQA(GPT), E5, and ALTER, showing the effectiveness of its complementary reasoning in a two-stage process. Appendix A.2 presents a comprehensive comparison of methods, including pre-trained models, fine-tuned architectures, and state-of-the-art LLM-based approaches. Our evaluation demonstrates that H-STAR consistently achieves superior performance across all benchmark metrics compared to all the baselines.
Performance Across LLMs.
@@ 3.2 Efficiency Analysis
Analysis. H-STAR employs a structured, step-by-step process: it first extracts query-relevant columns before selecting relevant rows. This progressive filtering minimizes token usage at each stage, thereby reducing computational overhead. As a result, H-STAR significantly reduces the average number of processed table cells—dropping from 159 to 18 on the WikiTQ dataset, from 88 to 13 on TabFact, and from 86 to 10 on FeTaQA. In contrast, alternative table manipulation methods, such as TabSQLify and DATER, yield considerably higher values. By effectively filtering out irrelevant information, H-STAR enables LLMs to focus on the right evidence for right reasoning, demonstrating superior efficiency across all datasets.